OpenStudy (khantahmina):

help! 3sin^2 x=cos^2(x)

OpenStudy (khantahmina):

@pooja195 @Mehek14

OpenStudy (khantahmina):

@TheSmartOne

TheSmartOne (thesmartone):

mhmm divide cos^2(x) on both sides Hint: \(\Large \frac{sin^2(x)}{cos^2(x)} = tan^2(x)\) and then divide by 3 on both sides, and you should be able to handle it from there mhmm

OpenStudy (khantahmina):

\[\tan^2(x)=(1/3)\]

OpenStudy (khantahmina):

n then do this?

OpenStudy (khantahmina):